A scuba diver, hired by an amusement park, collected $159 in nickels, dimes, and quarters at the bottom of a wishing well. There were 700 nickels, and 90 more quarters than dimes. How many quarters and dimes were thrown into the wishing well?
Answer by richwmiller(17219)

700 nickels is 700*.05=$35.00
$159-35=124
90*.25=$22.50
124-22.50=101.50
.25q+.1d=101.50
d=q
.35d=101.50
d=101.50/.35
d=290
q=290+90=380
n=700
380*.25=$95.00 in quarters
290*.10=$29.00 in dimes
700*.05=$35.00 in nickels